# The Challenge of Using LLMs to Simulate Human Behavior: A Causal Inference Perspective

### v2

@@ 6 Conclusion
This paper identifies a fundamental challenge in using LLMs to simulate experiments with human participants. Unlike human participants who exist independently of experiments, LLM-simulated individuals and environments are generated based on prompts. While LLMs’ ability to capture associations from training data is generally beneficial, it creates complications in experimental settings where changing one variable can unintentionally affect other unspecified variables meant to remain constant. This confounding makes it difficult to accurately simulate counterfactual human behavior. More generally and fundamentally, making the LLM blind to the experimental design leads to ambiguous prompts. Given LLMs are trained primarily on observational data, ambiguous prompts tend to favor correlational rather than causal observations, making variables unspecified in the prompt susceptible to the intended intervention. Because these challenges stem from ambiguous prompting strategies rather than model limitations, our theoretical framework suggests that it is not possible to fully resolve it through improved training data.
@@
We explored several potential solutions to this challenge. Adding detailed context to control for confounders showed promise but tends to introduce focalism which could invalidate results. A more promising approach consists in unblinded experimental designs that explicitly communicate randomization. This approach showed potential for reducing ambiguity, though its effectiveness may vary across models and contexts. Therefore, developing LLMs that can better follow the instruction to simulate counterfactual human behavior given unambiguous prompts can be a crucial direction for future research.
Our paper highlights a broader insight of using LLMs: rather than simply applying human subject protocols to LLMs, researchers need to develop new experimental practices tailored to LLMs’ unique properties. The field of LLM simulations is still emerging, and best practices remain to be established. Our work provides a theoretical foundation for understanding one of the key challenges and evaluating potential solutions. We hope this framework will guide future research in developing more reliable methods for leveraging LLMs in experimental research, while remaining mindful of their fundamental differences from human subjects.
@@ Appendix A Category, Product and Price Information
Note: The 40 categories were taken from DellaVigna and Gentzkow, (2019) and the top products and their regular prices from Walmart.com as of April, 2024. Due to the dynamic nature of retail pricing, actual prices may vary.
@@ Appendix B Simple prompts for eliciting correlation
Prompt 5: Correlation between price and competing price System: You, AI, are a customer. Your task is to fill in the blanks __. Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at: {price}. The price of a similar competing product from a different brand is           [a number with up to 2 decimal points]. Return example 1: XX.XX
Prompt 6: Correlation between price and expiration days System: You, AI, are a customer. Your task is to fill in the blanks __. Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at: {price}. The expiration date of the product is           [a whole number] days from now. Return example 1: 10
@@ Appendix C Covariance of unspecified variables
This section demonstrates how prevalent variables covary with each other. We consider a more detailed prompt C.1 that elicits demographic, contextual, and pre-treatment variables.
Prompt 7: Correlation between multiple variables System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: You are a consumer with the following characteristics: Age:   [a whole number] Gender:   Education level:   Household income:   [a whole number] Occupation:   Ethnicity:   Marital status:   Household size:   [a whole number] Number of children:   [a whole number] State of residence:   [state] Home ownership:   [e.g., “own,” “rent”] Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. You have purchased this product   [e.g., “frequently,” “occasionally,” “rarely”] in the past. The last time you saw this product, it was priced at \$  [a number with up to 2 decimal points]. You have   [e.g., “a lot of,” “some,” “limited”] storage space at home. Your monthly grocery budget is   [a whole number]. The expiration date of the product is   [a whole number] days from now. The product is currently priced at \${Price}. A similar competing product from a different brand is priced at \$  [a number with up to 2 decimal points]. Would you or would you not purchase the product?   [“purchase” or “not purchase”] Return example: 35, female, bachelor’s degree, 50000, software engineer, Asian, married, 3, 1, CA, own, occasionally, 2.99, some, 800, 60, 3.49, purchase
@@
Figure C.2 summarizes the covariance matrix between these variables. On one hand, this covariance matrix demonstrates that the LLM is capable of simulating many realistic correlations between covariates. For example, household income is positively correlated with budget. However, several covariates are correlated with the price change variable. Specifically, the focal price is correlated with product characteristics such as expiration days of the product, contextual variables such as competing prices, and pre-treatment variables such as purchase history. In an experiment with humans, these variables would either be held constant or be randomly distributed across treatment conditions in a way that is orthogonal to the treatment.
@@ Appendix D Simulation that controls for demographic variables
We use the following prompt template to first generate the distribution of demographic variables and then control for demographic variables. First, for each {category,product}, we leave the demographic variables as blank, and use the pinned version of gpt-4o-mini-2024-07-18 to generate 500 replications (personas) with a temperature of 1, giving us the joint distribution of demographic variables for that product. See Prompt D. To avoid imposing any prior assumption on Price, we leave {Price} as ”   [a number with up to 2 decimal points]” in this step. Then in the second step, we use the same prompt template but with the demographic variables filled in, i.e., all variables except the purchase decision filled in and the price being randomized. For each {category,product} and for each persona for that product, we vary the price, and elicit the purchase decision, while setting the temperature to 0, and then aggregate across the 500 personas to calculate the purchase probability. Then we aggregate across all products at the same relative price to construct the aggregate demand curve.
@@
Prompt 8: Control for demographic variables System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: You are a consumer with the following characteristics: Age:   [a whole number] Gender:   Education level:   Household income:   [a whole number] Occupation:   Ethnicity:   Marital status:   Household size:   [a whole number] Number of children:   [a whole number] State of residence:   [state] Home ownership:   [e.g., “own,” “rent”] Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at   [a number with up to 2 decimal points]. Would you or would you not purchase {product}?   [“purchase” or “not purchase”] Return example: 35, female, bachelor’s degree, 50000, software engineer, Asian,married, 3, 1, CA, own, 3.99, purchase
Appendix E Simulation that controls for demographic variables + competing price
@@
To provide a concrete example, we consider Coca-cola and asks what happens to the purchase probability when the price of a competing product is set at a certain level. We follow the exact same two-step procedure as Appendix D, except that we have fixed the price of a competing product to be the same as the regular price of the focal product when eliciting purchase decisions from the 500 personas.
@@
Prompt 9: Control for demographic variables+ competing price System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: You are a consumer with the following characteristics: Age: {age} Gender: {gender} Education level: {education} Household income: {income} Occupation: {occupation} Ethnicity: {ethnicity} Marital status: {marital status} Household size: {household size} Number of children: {number of children} State of residence: {state} Home ownership: {home ownership} Please consider the following product category: Soft Drinks - Carbonated. Suppose you are in a grocery store, and you see the following product in that category: Coca-Cola Soda Pop, 12 fl oz, 12 Pack Cans. The product is currently priced at \${Price}. A similar competing product from a different brand is priced at \$8.26. Would you or would you not purchase Coca-Cola Soda Pop, 12 fl oz, 12 Pack Cans?   [“purchase” or “not purchase”] Return example: purchase
@@
This mimics the scenario of an actual store, where the price of a competing product is set uniformly at a certain level. Different customers, each with their own characteristics, will come to the store and make their purchase decision. While it is plausible that some customers may notice the price of the competing product and use it as a reference point to make purchase decisions, not all customers are attentive and may be aware of the price of all other competing products. Therefore, it is implausible to assume that the purchase will be $100\%$ ($0\%$) when the price of the focal product is lower than or equal to (greater than) the price of competing products.
Appendix F Direct simulation of the experimental results using unblinded design
Prompt F is a prompt template we use to predict the experimental results using an unblinded design, where information about the experiment is directly provided. For each {category,product}, we replace {RegularPrice} with the corresponding price listed in Table A.1.
For Figure 7, we replace the {Prices} with the set of prices that we experiment with, from 0% to 200% of regular price. For example, when the regular price is \$4.68. {Prices} are ”\$0.00, \$0.94, \$1.87, \$2.81, \$3.74, \$4.68, \$5.62, \$6.55, \$7.49, \$8.42, \$9.36”. In this basic version, we leave Additional Instruction 1 and 2 as empty strings “”. We set the temperature to 0 and use ’gpt-4o-mini-2024-07-18’.
For Figure 7, in addition to considering the {Prices} to be from 0% to 200%, we also try a version from 0% to 100%, labeled as ”Lower_Half_Price_Range” and a version from 100% to 200%, labeled as ”Upper_Half_Price_Range”. It shows that the simulation result is sensitive to the range of prices provided in the experiment.
To assess whether this sensitivity is driven by another form of prompt ambiguity, where LLMs may make unintended assumptions on the context given the set of experimental prices, in Figure F.1 we provide more explicit instructions by replacing {Additional Instruction 1} with ”Note that some elements of the experiments are chosen arbitrarily by the researchers and you should not rely on them to make your prediction.” and replacing {Additional Instruction 2} with ”Note that the Set of Prices to be experimented is chosen arbitrarily and you should not make your prediction based on this Set of Prices.” However, we find that the results are still sensitive, suggesting that the issue are driven by the current generation of LLMs that we are testing being unable to closely follow such instruction, rather than prompt ambiguity. To assess the potential for future models to make consistent prediction, we further test ’o1-mini-2024-09-12’ and ’o1-preview-2024-09-12’. Because they do not allow for ’System’ and only accept the ’User’ part of the prompt, we concatenate the ’System’ and ’User’ together as our input prompt. Figure F.1 shows that when using ’gpt-4o-mini-2024-07-18’, the inconsistency still remains substantial despite explicit instruction not to rely too much on the experimental range to make the prediction. However, ’o1-mini-2024-09-12’ and ’o1-preview-2024-09-12’ demonstrates increasingly decreased inconsistencies.
[…]
Figure F.1: Unblinded design with explicit instruction to disregard experimental range: Multiple consecutive model versions.

### v3

@@ 6 Conclusion
This paper identifies a fundamental challenge in using LLMs to simulate experiments with human participants. Unlike human participants who exist independently of experiments, LLM-simulated individuals and environments are generated based on prompts. While LLMs’ ability to capture associations from training data is generally beneficial, it creates complications in experimental settings where changing one variable can unintentionally affect other unspecified variables meant to remain constant. This confounding makes it difficult to accurately simulate counterfactual human behavior.
Our exploration of potential solutions to this challenge reveals important insights for improving LLM simulations. While adding detailed context to control for confounders can help address the confounding issue, our analysis demonstrates that this approach can introduce focalism—causing simulated individuals to overweight explicitly mentioned variables relative to realistic human behavior. This trade-off between reducing confounding and maintaining ecological validity represents a fundamental consideration when designing LLM simulations with detailed prompts.
These empirical observations point to a deeper theoretical issue. Making the LLM blind to the experimental design leads to ambiguous prompts. Because these challenges stem from ambiguous prompting strategies rather than model limitations, our theoretical framework suggests that developing unambiguous prompting strategies constitutes a fundamental principle that complements other model improvement techniques.
Building on this theoretical understanding, our empirical findings demonstrate that unblinded experimental designs offer a particularly promising path forward. By explicitly communicating the randomization mechanism and experimental context, unblinded prompts reduce ambiguity and enable more accurate simulations. This approach shows consistent improvements across model generations and proves especially valuable when combined with fine-tuning, where it enhances the model’s ability to distinguish causal from correlational patterns in training data. The complementary nature of unblinding and fine-tuning suggests that developing LLMs capable of accurately simulating counterfactual human behavior when provided with clear, unambiguous prompts represents both a crucial and achievable direction for advancing the field.
@@
Our paper highlights a broader insight regarding the use of LLMs: rather than simply applying human subject protocols to LLMs, researchers must develop new experimental practices tailored to LLMs’ unique properties. The field of LLM simulations remains nascent, and best practices have yet to be established. Our work highlights the need for a new set of principles for designing LLM simulations and evaluation frameworks. We hope this will guide future research in developing more reliable methods for leveraging LLMs in experimental research while remaining mindful of their fundamental differences from human subjects.
Funding and Competing Interests Declarations
All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript. The authors have no funding to report.
@@ Appendix A Category, Product and Price Information
Note: The 40 categories were taken from dellavigna_uniform_2019 and the top products and their regular prices from Walmart.com as of April, 2024. Due to the dynamic nature of retail pricing, actual prices may vary.
@@ Appendix B Simple prompts for eliciting correlation
Prompt 7: Correlation between price and competing price System: You, AI, are a customer. Your task is to fill in the blanks $_$. Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at: {price}. The price of a similar competing product from a different brand is      [a number with up to 2 decimal points]. Return example 1: $X$.XX
Prompt 8: Correlation between price and expiration days System: You, AI, are a customer. Your task is to fill in the blanks $_$. Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at: {price}. The expiration date of the product is      [a whole number] days from now. Return example 1: 10
@@ Appendix C Covariance of unspecified variables
This section demonstrates how prevalent variables covary with each other, for which we use a more detailed prompt C.2 that elicits demographic, contextual, and pre-treatment variables. Figure C.1 summarizes the covariance matrix between these variables. On one hand, this covariance matrix demonstrates that the LLM is capable of simulating many realistic correlations between covariates. For example, household income is positively correlated with budget. However, several covariates are correlated with the price change variable. Specifically, the focal price is correlated with product characteristics such as expiration days of the product, contextual variables such as competing prices, and pre-treatment variables such as purchase history. In an experiment with humans, these variables would either be held constant or be randomly distributed across treatment conditions in a way that is orthogonal to the treatment.
Prompt 9: Correlation between multiple variables System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: You are a consumer with the following characteristics: Age:   [a whole number] Gender:   Education level:   Household income:   [a whole number] Occupation:   Ethnicity:   Marital status:   Household size:   [a whole number] Number of children:   [a whole number] State of residence:   [state] Home ownership:   [e.g., “own,” “rent”] Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. You have purchased this product   [e.g., “frequently,” “occasionally,” “rarely”] in the past. The last time you saw this product, it was priced at \$  [a number with up to 2 decimal points]. You have   [e.g., “a lot of,” “some,” “limited”] storage space at home. Your monthly grocery budget is   [a whole number]. The expiration date of the product is   [a whole number] days from now. The product is currently priced at \${Price}. A similar competing product from a different brand is priced at \$  [a number with up to 2 decimal points]. Would you or would you not purchase the product?   [“purchase” or “not purchase”] Return example: 35, female, bachelor’s degree, 50000, software engineer, Asian, married, 3, 1, CA, own, occasionally, 2.99, some, 800, 60, 3.49, purchase
@@ Appendix D Simulation that controls for demographic variables
We use Prompt D to first generate the distribution of demographic variables and then control for demographic variables. First, for each {category,product}, we leave the demographic variables as blank, and use the pinned version of gpt-4o-mini-2024-07-18 to generate 500 replications (personas) with a temperature of 1, giving us the joint distribution of demographic variables for that product. See Prompt D. To avoid imposing any prior assumption on Price, we leave {Price} as ”   [a number with up to 2 decimal points]” in this step. Then in the second step, we use the same prompt template but with the demographic variables filled in, i.e., all variables except the purchase decision filled in and the price being randomized. For each {category,product} and for each persona for that product, we vary the price, and elicit the purchase decision, while setting the temperature to 0, and then aggregate across the 500 personas to calculate the purchase probability. Then we aggregate across all products at the same relative price to construct the aggregate demand curve.
@@
Prompt 10: Control for demographic variables System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: You are a consumer with the following characteristics: Age:   [a whole number] Gender:   Education level:   Household income:   [a whole number] Occupation:   Ethnicity:   Marital status:   Household size:   [a whole number] Number of children:   [a whole number] State of residence:   [state] Home ownership:   [e.g., “own,” “rent”] Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at   [a number with up to 2 decimal points]. Would you or would you not purchase {product}?   [“purchase” or “not purchase”] Return example: 35, female, bachelor’s degree, 50000, software engineer, Asian,married, 3, 1, CA, own, 3.99, purchase
D.1 Simulation that also controls for competing price
@@
To illustrate the unintended effect of adding covariates in simulation, we focus on a product, Coca-cola, and asks what happens to the purchase probability when the price of a competing product is set at a certain level. We follow the exact same two-step procedure as Web Appendix D, except that we have fixed the price of a competing product to be the same as the regular price of the focal product when eliciting purchase decisions from the 500 personas.
@@
Prompt 11: Control for demographic variables+ competing price System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: You are a consumer with the following characteristics: Age: {age} Gender: {gender} Education level: {education} Household income: {income} Occupation: {occupation} Ethnicity: {ethnicity} Marital status: {marital status} Household size: {household size} Number of children: {number of children} State of residence: {state} Home ownership: {home ownership} Please consider the following product category: Soft Drinks - Carbonated. Suppose you are in a grocery store, and you see the following product in that category: Coca-Cola Soda Pop, 12 fl oz, 12 Pack Cans. The product is currently priced at \${Price}. A similar competing product from a different brand is priced at \$8.26. Would you or would you not purchase Coca-Cola Soda Pop, 12 fl oz, 12 Pack Cans?   [“purchase” or “not purchase”] Return example: purchase
@@
This mimics the scenario of an actual store, where the price of a competing product is set uniformly at a certain level. Different customers, each with their own characteristics, will come to the store and make their purchase decision. While it is plausible that some customers may notice the price of the competing product and use it as a reference point to make purchase decisions, our step-wise non-smooth demand curve is implausible in real-world scenarios where not all customers are attentive or are aware of the price of all other competing products.
Appendix E Non-monotonic impact of controlling for covariates
Table E.1 gives an overview of the variables used and the corresponding results documented in Section 3.
Table E.1: Sequential Covariate Addition in Sensitivity Analysis
Chunks Total New New Covariates MAE
Covariates Covariates (Names)
1 14 14 Age, US Citizenship, Education Level, Employmen… 0.2250
2 15 1 Tightwad-Spendthrift Score 0.0971
3 17 2 Discount Rate, Present Bias 0.0989
4 18 1 Risk Aversion 0.1413
5 19 1 Loss Aversion 0.1478
6 20 1 Financial Literacy 0.1524
7 21 1 Numeracy 0.1714
8 22 1 score mentalaccounting 0.1646
9 23 1 score maximization 0.1700
10 24 1 score minimalism 0.1842
11 25 1 score GREEN 0.2025
12 30 5 score agreeableness, score extraversion, score … 0.1929
  • Note: This table presents the sequential addition of covariates in the sensitivity analysis. Stage 1 includes sociodemographic characteristics (age, citizen status, education, employment status, household size, marriage, political affiliation, political views, race, region, religion, religious attendance, sex, and total family income). Stages 2-11 add individual behavioral economics and psychological measures. Stage 12 adds Big Five personality traits (agreeableness, extraversion, neuroticism, openness, conscientiousness). Mean Absolute Error (MAE) represents prediction error at each stage.
Prompt E shows the final template that includes all the covariates we tested at Stage 12. At stage 1, only the demographic variables (the first 14 items) were included. Each subsequent stage systematically added behavioral and psychological measures: tightwad-spendthrift (stage 2), temporal discounting (stage 3), risk aversion (stage 4), loss aversion (stage 5), financial literacy (stage 6), numeracy (stage 7), mental accounting (stage 8), maximization (stage 9), minimalism (stage 10), environmental attitudes (stage 11), and finally the Big Five personality traits (stage 12).
Prompt 12: Prompt Templates Based on Detailed Covariates: Part I System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text.
User: You are a consumer with the following characteristics: # Demographics:
- Geographic region: {region}
- Gender: {sex}
- Age: {age}
- Education level: {education}
- Race: {race}
- Citizen of the US: {citizen_status}
- Marital status: {marriage}
- Religion: {religion}
- Religious attendance: {religious_attendance}
- Political affiliation: {political_affiliation}
- Income: {total_family_income}
- Political views: {political_views}
- Household size: {household_size}
- Employment status: {employment_status}
# Tightwad-Spendthrift: {score_ST-TW} ({pct_spendthrift} percentile)
<<note: The score ranges from 4 to 26. Lower scores (4-11) indicate difficulty spending money, while higher scores (19-26) indicate difficulty controlling spending.>>
# Discount, Present Bias:
<<note: These are implied rates computed from your time-value of money preferences. Higher values of the discount rate imply greater impatience. Higher values of present bias imply greater departure from normative economic behavior.>>
- Discount: {score_discount} ({pct_discount} percentile)
- Present Bias: {score_presentbias} ({pct_presentbias} percentile)
# Risk aversion: {score_riskaversion} ({pct_riskaversion} percentile)
<<note: Higher scores indicate a greater tendency for risk aversion in a choice between a sure-amount and lottery payout.>>
Prompt 12: Prompt Templates Based on Detailed Covariates: Part II # Loss aversion: {score_lossaversion} ({pct_lossaversion} percentile)
<<note: Higher scores indicate a greater tendency for loss aversion in a choice between a sure-amount and a lottery payout.>>
# Financial Literacy: {score_finliteracy} ({pct_finliteracy} percentile)
<<note: The score ranges from 0 to 8, and a higher score indicates you correctly answered more questions related to general financial literacy.>>
# Numeracy: {score_numeracy} ({pct_numeracy} percentile)
<<note: The score ranges from 0 to 8, and a higher score indicates you correctly answered more questions related to numeracy.>>
# Mental Accounting: {score_mentalaccounting} ({pct_mentalaccounting} percentile)
<<note: The score ranges from 0 to 100 percent, and higher scores indicate a greater adherence to the principles of mental accounting proposed by Thaler: segregate gains, integrate losses, segregate a small gain from a large loss, and integrate a small loss with a large gain.>>
# Maximization: {score_maximization} ({pct_maximization} percentile)
<<note: The score ranges from 1 to 5, and higher scores indicate a tendency to optimize rather than satisfice when making decisions.>>
# Minimalism: {score_minimalism} ({pct_minimalism} percentile)
<<note: The score ranges from 1 to 5, and a higher score indicates a higher preference for minimalism.>>
# GREEN: {score_GREEN} ({pct_green} percentile)
<<note: The score ranges from 1 to 5, and higher scores indicate a higher affinity for environmentalism.>>
# Big 5 Personality:
<<note: Openness reflects curiosity and receptiveness to new experiences, Conscientiousness indicates self-discipline and goal-directed behavior, Extraversion measures sociability and assertiveness, Agreeableness reflects compassion and cooperativeness, and Neuroticism captures emotional instability and susceptibility to negative emotions. Each score ranges from 1 to 5, and a higher score indicates a greater display of the associated traits.>>
- Extraversion: {score_extraversion} ({pct_extraversion} percentile)
- Agreeableness: {score_agreeableness} ({pct_agreeableness} percentile)
- Conscientiousness: {wave1_score_conscientiousness} ({pct_conscientiousness} percentile)
- Openness: {score_openness} ({pct_openness} percentile)
- Neuroticism: {score_neuroticism} ({pct_neuroticism} percentile)
Appendix F Theoretical Framework
F.1 Impossibility of Ambiguous Prompting
Theorem F.1.
Proof.
F.2 Illustration of the underlying DGP that LLM simulation mimics if prompts are interpreted differently
$Y=\beta_{M}D+U$
where $\beta_{M}$ represents the causal effect of $D$ on $Y$ through a set of realistic mechanism $M$, and $U$ represents unobserved factors. In a well-designed experiment, we have $\text{Cov}(U,D)=0$, making it possible to identify $\beta_{M}$ and estimate $P(Y|do(D))$ from data on $(Y,D)$ alone. But under a blinded design with an ambiguous prompt, the LLM does not know it should mimic an experimental environment. That is, the LLM does not have enough information to draw $U$ from the appropriate distribution (literally or figuratively - we do not attempt to model how the LLM makes inference but rather present a conceptual model). If the model is mostly trained on observational data, it is mostly likely going to mimic observational data patterns, where $\text{Cov}(U,D)\neq 0$. As a result, the LLM’s responses will reflect confounded relationships rather than the intended causal effects. Figure F.1 summarizes two possible DGPs mimicked by the LLM under this simple ambiguous prompt.
DDMMYYUU(a) DGP for Experimental DataDDMMYYUU(b) DGP for Observational Data
Figure F.1: Possible DGPs mimicked by LLM simulations under ambiguous simple prompts
As discussed in Section 3, one might attempt to address this issue by adding more variables $X$ to the prompt. In principle, if $\text{Cov}(U,D|X)=0$, then controlling for $X$ can help restore identification. However, simply specifying covariates in the LLM prompt does more than condition on them statistically. Including $X$ may cause the simulation to alter the underlying decision-making mechanisms $M$ and shift toward a new set of mechanisms $M^{\prime}$ and produce an environment that differs substantially from the original scenario – the estimated parameter $\beta_{M^{\prime}}$ may have limited ecological validity, since it no longer reflects the realistic mechanisms $M$ that were originally of interest (as we illustrated in Figure 4). Figure F.2 summarizes two possible DGPs mimicked by the LLM under this more complicated ambiguous prompt.
DDMMYYUUXX(a) Adding $X$ may address confoundingDDMMYYUUXX(b) Adding $X$ may also sacrifice ecological validity
Figure F.2: Potential DGPs mimicked by LLM simulations when controlling for additional variables
F.3 Implications for Model Development
The impossibility theorem extends to fine-tuning and RAG approaches. Consider experimental data where Coca-Cola prices were randomized while Pepsi prices remained fixed. Fine-tuning on this data with ambiguous prompts cannot guarantee the model learns the causal relationship $\beta_{M}$. The model may learn the regression coefficient $\hat{b}=\beta_{M}+\text{Cov}(U,D)/\text{Var}(D)$ while maintaining a prior that $\text{Cov}(U,D)\neq 0$, leading to biased causal inference.
Moreover, even if the model correctly learns independence assumptions for the training scenario ($q_{1}$), it may fail to generalize appropriately to new scenarios ($q_{2}$). When asked about cases where Pepsi prices vary and Coca-Cola prices are fixed, the model might incorrectly maintain the independence assumption from training, when correlation with unobserved factors would actually exist. Therefore, while fine-tuning and RAG remain valuable for incorporating domain knowledge, they are complements to rather than substitutes for unambiguous prompt designs.
Appendix G Fine-tuning with additional irrelevant data
To test the robustness of different fine-tuning strategy when additional dataset is included, we sample $10,000$ observations from the Amazon dataset berke2024open and incorporate them into fine tuning using Prompt G).
Prompt 13: Prompt for Amazon data fine-tuning System: You, AI, are a customer. Your task is to fill in the blank ___. Return the completed information without extra text. User: Consider the product category: {category}. Suppose you are shopping online, and you see the product {product}. The product is priced at \${price}. You decide to ___ [“purchase” or “not purchase”]. Return example 1: purchase Return example 2: not purchase